OpenStudy (anonymous):

x2 – 16x + 15 = 0 a) x = 3, x = 5 b) x = 1, x = 15 c) x = -3, x = -5 d) x = -1, x = -15

OpenStudy (aroub):

b) x = 1, x = 15

OpenStudy (anonymous):

factor: (x-15)(x-1)=0 x-15=0 x=15 x-1=0 x=1

OpenStudy (anonymous):

so its B